# Charge transport and mode transition in dual-energy electron beam diodes

### v1

@@ Abstract
The charge transport mode transition when the two electron beams have nearly identical energies (e.g., $k_{1}=\beta_{1}/\beta_{2}=0.95$) is shown in Fig. 3(d). Due to the negligible difference between the SCL currents of the two electron beams ($\text{e}_{1}$ and $\text{e}_{2}$), the system exhibits behavior analogous to that of a single electron beam in terms of current transmission. That is, two groups of electrons can be treated as the same electrons while determining the transmitted current. When the total injected current $J_{0}$ reaches the SCL current, the transport for the two groups of electrons can be simultaneously suppressed with partial reflection [M5, also see Fig. 2(a5)]. However, the low-energy electrons can be fully reflected when the total current is sufficiently large [see Fig. 3(d)]. Figure 3(h) presents the charge transport mode transition pathway ($\text{M1}\rightarrow\text{M2}\rightarrow\text{M5}\rightarrow\text{M4}$) for $k_{1}\rightarrow 1$, which differs from the observed pathway for $k_{1}<1$ [see Fig. 2(e)].
@@
Figure 4: Transmitted current density $J_{\text{tran}}$ versus the total injected current density $J_{0}$. The prediction from the proposed formula [Eq. (3)] agrees well with PIC simulation results with the electron beam number $n=3$.
@@
$J_{\text{tran}}=\left\{\begin{aligned} &\frac{n+1-i}{n}J_{0},\ J_{0}\in\left[\frac{(n+2-i)nJ_{\text{SCL},i-1}}{n+1-i},\ nJ_{\text{SCL},i}\right),\\ &(n+1-i)J_{\text{SCL},i},\ J_{0}\in\left[nJ_{\text{SCL},i},\ \frac{(n+1-i)nJ_{\text{SCL},i}}{n-i}\right),\end{aligned}\right$. (3)
where $i$ varies from 1 to $n$. The maximum transmitted current $J_{\text{tran}}^{\text{max}}=J_{\text{SCL},n}$, independent of the injected current $J_{0}$, is reached when the highest-energy electron beam $\text{e}_{n}$ reaches its SCL current, i.e., $J_{n}=J_{\text{SCL},n}$.
Conclusion—In conclusion, we explicitly determined five complete modes for charge transport in dual-energy electron beam driven diodes, which are modulated by controlling the two electron beam parameters. The initial velocity ratio and the injected current ratio are identified as key factors governing parameters; the former determines the charge transport modes and their transitions, while the latter modulates the characteristics of the transmitted current. These discoveries provide a fundamental advance in multi-energy beam physics, enabling precise tuning strategies for electron beam driven diodes, especially for applications in modern vacuum electronic devices.

### v2

@@ Abstract
The charge transport mode transition when the two electron beams have nearly identical energies (e.g., $\beta_{1}/\beta_{2}=0.95$) is shown in Fig. 3(d). Due to the negligible difference between the SCL currents of the two electron beams ($\text{e}_{1}$ and $\text{e}_{2}$), the system exhibits behavior analogous to that of a single electron beam in terms of current transmission. That is, two groups of electrons can be treated as the same electrons while determining the transmitted current. When the total injected current $J_{0}$ reaches the SCL current, the transport for the two groups of electrons can be simultaneously suppressed with partial reflection [M5, also see Fig. 2(a5)]. However, the low-energy electrons can be fully reflected when the total current is sufficiently large [see Fig. 3(d)]. Figure 3(h) presents the charge transport mode transition pathway ($\text{M1}\rightarrow\text{M2}\rightarrow\text{M5}\rightarrow\text{M4}$) for $\beta_{1}/\beta_{2}\rightarrow 1$, which differs from the observed pathway for $\beta_{1}/\beta_{2}<1$ [see Fig. 2(e)].
@@
Figure 4: Transmitted current density $J_{\text{tran}}$ versus the total injected current density $J_{0}$. The prediction from the proposed theoretical formula [Eq. (3)] agrees well with PIC simulation results, for example, with the electron beam number $n=3$.
@@
$J_{\text{tran}}(J_{0})=\left\{\begin{aligned} &\frac{n+1-i}{n}J_{0},\ J_{0}\in\left[J_{\text{a}},\ J_{\text{b}}\right),\\ &(n+1-i)J_{\text{SCL},i},\ J_{0}\in\left[J_{\text{b}},\ J_{\text{c}}\right),\end{aligned}\right$. (3)
where $J_{\text{a}}={(n+2-i)nJ_{\text{SCL},i-1}}/{(n+1-i)}$, $J_{\text{b}}=nJ_{\text{SCL},i}$, $J_{\text{c}}={(n+1-i)nJ_{\text{SCL},i}}/{(n-i)}$, and $i$ varies from 1 to $n$. The maximum transmitted current $J_{\text{tran}}^{\text{max}}=J_{\text{SCL},n}$, independent of the injected current $J_{0}$, is reached when the highest-energy electron beam $\text{e}_{n}$ reaches its SCL current, i.e., $J_{n}=J_{\text{SCL},n}$.
Conclusion—In conclusion, we explicitly determined five complete modes for charge transport in dual-energy electron beam driven diodes, which are modulated by controlling the two electron beam parameters. The initial velocity ratio and the injected current ratio are identified as governing parameters; the former determines the charge transport modes and their transitions, while the latter modulates the characteristics of the transmitted current. The discovery provides a fundamental advance in multi-energy beam physics, enabling precise tuning strategies for electron beam driven diodes, especially for applications in modern vacuum electronic devices.